4) Suppose there is a 3-year bond with a $1000 face value, 30% annual coupon payments and a 20% annual yield to maturity. a) Without any calculation, briefly explain whether this bond will be selling a premium or a discount b) Calculate the price of this bond. c Calculate the duration of this bond. d) Suppose the interest rates in the economy rise by 5 percentage points immediately after someone bought this bond. Show a calculation using duration for what should happen to the price of this bond immediately. Suppose now one year passes and the interest rates in the economy are still 5 percentage points higher. If the person that owns this bond sells it for fair value, what would be their one year rate of return? Show a calculation. e) 5 For each of the following state whether it is True or False and provide a brief explanation (I or 2 sentences) as to why it is true or false. a) b) c) All else the same, a higher coupon rate will mean a higher price for a standard coupon bond. All else the same, a higher coupon rate will mean a higher duration for a standard coupon bond. All else the same, a shortage in the money market will lead to the price of bonds falling.
Answer #1

a) since here the coupon payment is higher than yield to maturity the bond is selling at premiun
b)price of bond can be found uisng below formulae
Price=(coupon*(1-((1+i)^-n))/i)+(issue price*(1+i)^-n)
Coupon=face value*coupon rate
=1000*30%=300
i=20%
n=3 years
issue price=1000
substituting in formuale we get it as 1210.65
c)duration of bond below

Inputs Rate Convention: 1 EAR, 0 APR Annual Coupon Rate (CR) Yield to Maturity (Annualized) (y) Number of Pavments/ Year (NOP

BOND DURATION Basics 3 Inputs 4 Rate Convention 1 EAR, 0 APR 5 Annual Coupon Rate (CR) 6 Yield to Maturity (Annualized) (y) 7
Maculay duration 2.41 years

d)the price of bond will come down since here the interest rate increased. Since the price and yield are inversely proportional to each other
Bond Price Change=(Duration×Yield Change*100)
=-2.41*5%=-12.05%

e)let us find price of bond after one year
Price=(coupon*(1-((1+i)^-n))/i)+(issue price*(1+i)^-n)
Coupon=face value*coupon rate
=1000*30%=300
i=25%
n=2 years
issue price=1000
substituting in formuale we get it as 1072
return=(final/initial)-1
=(1072/1210.65)-1=-11.45%